In a manufacturing process, a transparent film is being bonded to a substrate as shown in the sketch. To cure the bond at a temperature "To", a radiant source is used to provide a heat flux, all of which is absorbed at the bonded surface. The back of the substrate is maintained at T1 while the free surface of the film is exposed to air at a certain temperature shown below with a corresponding heat transfer coefficient h. Given the values below, determine the equivalent resistance because of the thickness of the substrate (answer in m2-K/W)


1
Expert's answer
2021-04-12T01:54:56-0400

Even if the diagram was not included, we can consider the diagram below



The thermal circuit is shown in the Figure below



Equivalent resistance= "R_{Conv}+R_{Conv, f}+R_{Conv, s}"

"\frac{1}{hA}+\frac{L_f}{k_fA}+\frac{L_s}{k_sA}"

"A(\frac{1}{50A}+\frac{0.25 \times0.001}{0.025A}+\frac{0.001}{50A})=0.03002 m^2K/W"